An aircraft flying a heading a 60° at air speed 200 ms-1 moves relative to the ground due east (909) at 225 ms-1, If the plane is flying at constant pressure, what is its rate of change in altitude assuming a steady pressure field and f 104 81? 3.
Answer #1

Given : vair = 200 m/s , vground = 225 m/s , f = 10-4 s-1

Solution:

Visual of given situation :

Nain 60 30、 oC

Using cosine rules to find geostrophic wind speed :

vg2 = vair2 + vground2 -2(vair)(vground)cos(30)

= (200)2+(225)2-2(200)(225)cos(30)

= 12682.71

vg = sqrt12682.71 = 112.6 m/s

Geopotential height gradient can be determined as :

0 V)

rac{partial z}{partial n} = rac{fv_g}{g}

(10-4) (112.6) 9.81

= 1.148 × 10-3 mm-1

Applying sine rule to get the angle (alpha) :

air sina sin(30)

1200sin (30) 112.6

= 62.6°

The component of geopotential height gradient in the direcrion of

vground is given by:

9062.6) On = -1.0192 ×10-3 mm-1

The rate of change of altitude is given by:

-1.0192×10-3* vground

= -1.0192×10-3×225

= -0.229 ms-1

Answer : -0.229 ms-1